Ronald has a 12% solution of the fertilizer Super Grow. How much pure super grow should he add to the mixture to get 32oz of a 17.5% concentration
Answer by ptaylor(2198)

Let x=amount of pure Super Grow that needs to be added
Then 32-x=amount of 12% solution

Now we know that the amount of pure Super Grow in the 12% solution (0.12(32-x))plus the amount of pure Super Grow added (x) has to equal the amount of pure Super Grow in the final mixture (0.175(32)). So our equation to solve is:
0.12(32-x)+x=0.175*32 get rid of parens
3.84-0.12x+x=5.6 subtract 3.84 from both sides
3.84-3.84-0.12x+x=5.6-3.84 collect like terms
0.88x=1.76 divide both sides by 0.88
x=2oz----------------amount of pure Super Grow that needs to be added
CK
0.12*30+2=0.175*32
3.6+2=5.6
5.6=5.6

Hope this helps---ptaylor